Child pages
  • 201610 - Public report
Skip to end of metadata
Go to start of metadata

You are viewing an old version of this page. View the current version.

Compare with Current View Page History

« Previous Version 4 Current »


  • October 24th: The red link to GSIX was currently down and also the link to gu-br2 went down because of a fiber problem which is fixed now.
  • October 13th: A broken FPC was replaced in hb-r2
  • October 11th: BGP IPv4 sessions between NORDUnet and SUNET went down due to misconfiguration in NORDUnet NCS
  • October 7th: The green link to GSIX was down for few minutes. Customer did not have any maintenance. No RFO found.
  • October 4th: Problem in our optical network between hv-r2 and gbg7-r2.


  • October 30th: Supplier performed replacement of fiber cable in Gävle on the red network.
  • October 26th: A service that is going between tug-r1 and mdh-r2 was recreated on our optical NMS.
  • October 16th: Fiber supplier performed repair/replacement of fiber cable, Morjärv-Kälvudden.
  • October 14th: A maintenance was carried out in one of the MDH routers
  • October 12th: We performed a planned storage maintenance on the Adobe Connect system for SUNET.
  • October 11th: A broken FPC card needs to be replaced in router in Göteborg.
  • October 7th: A new service was created between core routers in GBG7 and BOS
  • October 6th: SUNET relocated tne FPC card to new SUNET-C router.
  • October 3rd: Fiber optimization work was carried out in Fredhäll site on SUNET-C network.






1 pcs 00:32

1 pcs 01:01

2 pcs 01:33


1 pcs 00:04

3 pcs 06:06

4 pcs 06:10


3 pcs 15:22

5 pcs 1d 02:10

8 pcs 1d 17:32


0 pcs

0 pcs

0 pcs


0 pcs

0 pcs

0 pcs

  • No labels